OpenStudy (anonymous):

Given the function f(x)=2x-1 sketch the following: (a) |f(x)| (B) 2f(x) (C) f(x)+2 (D) f(x-3) How do i do this? I am thinking for the f(x) function, to place it into each graph (a)-(d) then solve and finally graph the functions of each of my answers.....But i am sort of stuck with evaluationg.. E.g (c) 2x-1 +2 = 2x+1 OR 4x+1?? A bit confused...help please..

OpenStudy (calculusfunctions):

First graph the given function.

OpenStudy (anonymous):

Ok thanks but when i do...how would (a)-(b) from the function graph.

OpenStudy (anonymous):

dont forget for (a) it is taking about the absolute value

OpenStudy (anonymous):

I am thinking of evaluating the f(x) into each but the evaluating is giving me sort of a problem. Can sombody try one? I will use that as guide...

OpenStudy (anonymous):

Hi Farrah, it says given the function f(x) = 2x -1 what is missing here is a range of vaues for x in which to sketch, so I take it you can chose your own, you might try zero to 10 for example.

OpenStudy (anonymous):

I think i get what your saying : thanks BPD... For the f(x) i find the ranges of values x then use that to graph. (a)--(d) ?

OpenStudy (anonymous):

I think you should be given the range for x in orer to sketch it.(BTW sketch is not the same as plot) so you ust need a few values of x. Remember absolute vale !f(x)! can ony be positive.

OpenStudy (anonymous):

2f(x) will ust double the value of f(x)

OpenStudy (anonymous):

Once you have the values of f(x) for the given values of y you ust add 2 to find f(x)+2, but for d, again you will need to work this out e.g. if x = 1, f (x -3) = f (1-3) = f(-2) = 2(-2) -1 = -5

OpenStudy (anonymous):

Okay thank you...that was most helpful